OpenStudy (anonymous):

PLEASE HELP WILL GIVE MEDAL AND WILL FAN The following function represents the profit P(n), in dollars, that a concert promoter makes by selling tickets for n dollars each: P(n) = -250n^2 + 3,250n - 9,000 Part A: What are the zeroes of the above function, and what do they represent? Show your work. (4 points) Part B: Find the maximum profit by completing the square of the function P(n). Show the steps of your work. (4 points) Part C: What is the axis of symmetry of the function P(n)? (2 points)

OpenStudy (anonymous):

@jdoe0001 @pgpilot326

OpenStudy (jdoe0001):

hmm have you done quadratic factoring I assume? like say factoring \(\bf x^2+4x+4\to (\qquad ?)(\qquad? )\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(x+2)(x+2)

OpenStudy (jdoe0001):

right

OpenStudy (jdoe0001):

so part A is just asking for the zeros of the function or "solution" or "roots" which you can get by just setting P(n) to 0, that is \(\bf P(n) = -250n^2 + 3,250n - 9,000\implies 0= -250n^2 + 3,250n - 9,000 \\ \quad \\ simplified\implies 0=-250(n^2+13n-36)\implies 0=n^2+13n-36\)

OpenStudy (jdoe0001):

and you can get those by just factoring the quadratic

OpenStudy (jdoe0001):

hmm actually wait a sec

OpenStudy (jdoe0001):

I took out the -250 as common factor... need to change the signs

OpenStudy (jdoe0001):

\(\bf P(n) = -250n^2 + 3,250n - 9,000\to 0= -250n^2 + 3,250n - 9,000 \\ \quad \\ simplified\implies 0=-250(n^2-13n+36)\implies 0=n^2-13n+36\) anyhow, you can also factor that

OpenStudy (anonymous):

when i factored it i got (n−9)(n−4) so would the 0s be -9 and -4?

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

welll hold .. not quite... the roots are n-9 and n-4 yes the value for "n" would be \(\large 0=(n-9)(n-4)\to \begin{cases} 0=n-9\to &9=n \\ \quad \\ 0=n-4\to &4=n \end{cases}\)

OpenStudy (jdoe0001):

you know what a "perfect square trinomial" is, right?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

for part B \(\bf 0=n^2-13n+36\implies 0=(n^2-13n)+36 \\ \quad \\ 0=(n^2-13n+{\color{red}{ \square }}^2)+36\) any ideas what our missing number there is, to get a "perfect square trinomial" ?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

i tried finding the maximum and i got 6.5 is that correct?

OpenStudy (jdoe0001):

well... let us notice that the middle term of a perfect square trinomial is 2 * left term * right term so one could say that yeap... 6.5 is correct \(\bf 13x=2\cdot x\cdot \square \implies \cfrac{13\cancel{ x }}{2\cancel{ x }}=\square \)

OpenStudy (anonymous):

Would the axis of symmetry be 250?

OpenStudy (jdoe0001):

well dunno yet on the axis, that'd depend on part B though

OpenStudy (anonymous):

wait isn't part b done already because we determined that the maximum is 6.5 and thats what part b is asking for?

OpenStudy (jdoe0001):

well.. not quite

OpenStudy (jdoe0001):

so, keep in mind that, all we're doing is borrowing from our good friend Mr Zero, 0 so if we ADD 13/2, we have to also SUBTRACT 13/2 thus \(\bf 0=n^2-13n+36\implies 0=(n^2-13n)+36 \\ \quad \\ 0=\left[n^2-13n+\left({\color{red}{ \frac{13}{2} }}\right)^2\right]+36-\left({\color{red}{ \frac{13}{2} }}\right)^2 \\ \quad \\ 0=\left[n-\left(\frac{13}{2}\right)\right]^2+36-\frac{13}{2}\implies 0=\left[n-\left(\frac{13}{2}\right)\right]^2+\frac{59}{2}\)

OpenStudy (anonymous):

?

OpenStudy (jdoe0001):

heheh

OpenStudy (jdoe0001):

that's what part B was asking, to "complete the square" so you can find the maximum profit

OpenStudy (anonymous):

so whats the answer for part b? 0=[n−(13/2)]^2+59/2

OpenStudy (jdoe0001):

so the vertex of that parabola is at \(\bf 0=\left[n-{\color{blue}{ \left(\frac{13}{2}\right)}}\right]^2{\color{blue}{ +\frac{59}{2}}}\qquad vertex \ \left({\color{blue}{ \frac{13}{2}, \frac{59}{2}}}\right)\)

OpenStudy (anonymous):

so the vertex is (6.5,29.5)

OpenStudy (jdoe0001):

recall that before simplifying, it had a negative leading term coefficient that means the parabola is going downwards so yes, let us use (6.5, 29.5)

OpenStudy (anonymous):

which means that the axis of symmetry is 6.5

OpenStudy (jdoe0001):

 yeap, the axis of symmetry is \(\bf x = 6.5\)